OpenStudy (anonymous):

A scale drawing of a building has the scale 0.5 in. 6 ft a wall is 30 ft long. How long will the wall be on the drawing.

OpenStudy (anonymous):

can you clarify that problem is confusing to read

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

so .5 inches = 1 foot?

OpenStudy (anonymous):

and you want to know how long a 30 foot wall and a 6 foot wall would be if drawn to the scale .5 in = 1 foot. am i understanding you correctly?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

first you set up a ration of .5 in = 1 foot ( 12 in) therefore .5 in = 12 in. So the 30 foot wall has 360 inches in real life so you set .5 in = 12 inches and x in = 360 in and solve for x and you get x=15 which means that in your drawing a 30 foot wall will be represented by a 15 in line

OpenStudy (anonymous):

You solve the other wall of 6 feet lenght by doing the same steps

OpenStudy (anonymous):

can you just write all that in numbers because I can't understand words.

OpenStudy (ajprincess):

I guess it is this way. 6 feet of real building is represented by 0.5 inches in the drawing. 1 foot = 12 inches 6 feet = 6*12=72inches 30 feet = 30*12=360 inches Let us assume that 30 feet is represented by x inches in the drawing. \[0.5:72=\frac{0.5}{72}\] \[x:360=\frac{x}{360}\] \[\frac{x}{360}=\frac{0.5}{72}\] Find x. Does that help? @Maddie1234

OpenStudy (radar):

Is it really necessary to convert the wall dimensions to inches. When the scale is stated as .5 in = 6 ft. Then it is given that a wall is 30 ft long. You need only to find how many 6 ft increments are contained in 30 ft. That would be 5. So now you only need 5 times the .5 in. So the drawing will need to only be 5(.5) or 2.5 (2 1/2 in.) long.

OpenStudy (radar):

To say it with numbers then, let x = the required length on the drawing: x (.in) =(30/6) (.5 in)

OpenStudy (radar):

By the way, there is nothing wrong with ajprincess's approach, if you desire to convert all the building dimensions to inches.